Mechanics

The displacement of a body is represented by the equation X = kt^2 draw v t graph and a t graph

DreamKiller007 Gaming
6 Years agoGrade

1 Answer

Vikas TU
ApprovedApproved Tutor Answer6 Years ago
Dear student 
The displacement = kt^2 
Velocity = dx / dt 
V = 2kt 
So it will be a straight line with slope = 2k 
Acceleration = dv / dt 
So , a = dv /dt 
a = 2k 
So , acceleration will be constant with time.